OpenStudy (anonymous):

Solve the following system of equations please? :) My question got deleted before I could see the answers you guys gave! x - y = 19 2x + y = 2

OpenStudy (anonymous):

how'd you get that? i have to show my work. :)

OpenStudy (anonymous):

WHOA, that was wrong, let me re-do with work.

OpenStudy (anonymous):

thankyou <3

OpenStudy (anonymous):

Okay, I got the answer. x=7 y=-12 Now I'll show the working.

OpenStudy (anonymous):

Using the substitution method, rearrange the first equation to "x=" Add y to the other side and you get x=19+y Now that we know what x is, plug it into the second equation for the x value. 2(19+y)+y=2 Solve as you would a normal problem. 38+2y+y=2 38+3y=2 3y=-36 y=-12. Now plug in -12 for y in any other the two equations. 2x+(-12)=2 2x=14 x=7 or x-(-12)=19 x+12=19 x=7

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

Got it? :)

OpenStudy (anonymous):

might be easier just to add them and get \[3x=21\] so \[x=7\]